### v1

# Optimal Equivariant Matchings on the 6-Cube With an Application to the King Wen Sequence

Abstract
We characterize perfect matchings on the Boolean hypercube $\{0,1\}^{n}$ that are equivariant under the Klein four-group $K_{4}=\langle\mathrm{comp},\mathrm{rev}\rangle$ generated by bitwise complement and reversal. For $n=6$, we prove there exists a unique $K_{4}$-equivariant matching minimizing total Hamming cost among matchings using only $\mathrm{comp}$ or $\mathrm{rev}$ pairings, achieving cost 120 versus 192 for the complement-only matching. The optimal matching is determined by a simple “reverse-priority rule”: pair each element with its reversal unless it is a palindrome, in which case pair with its complement. We verify that the historically significant King Wen sequence of the I Ching is isomorphic to this optimal matching. Notably, allowing $\mathrm{comp}\circ\mathrm{rev}$ pairings yields lower cost (96), but the King Wen sequence follows the structurally simpler rule. All results are formally verified in Lean 4 with the Mathlib library.
1 Introduction
The Boolean hypercube $\{0,1\}^{n}=\{0,1\}^{n}$ admits a natural action by the Klein four-group $K_{4}\cong\mathbb{Z}_{2}\times\mathbb{Z}_{2}$, generated by bitwise complement $\mathrm{comp}$ and bit reversal $\mathrm{rev}$. A $K_{4}$-equivariant perfect matching is a partition of $\{0,1\}^{n}$ into pairs such that paired elements lie in the same $K_{4}$-orbit.
This paper addresses the following optimization problem:
Among $K_{4}$-equivariant perfect matchings on $\{0,1\}^{n}$ where each pair uses either $\mathrm{comp}$ or $\mathrm{rev}$, which minimize total Hamming distance (Definition 2.3)?
For $n=6$, we prove uniqueness: there is exactly one optimal matching of this form, determined by a simple greedy rule. As an application, we verify that the King Wen sequence—a traditional ordering of the 64 hexagrams of the I Ching into 32 pairs—is isomorphic to this unique optimal matching. We also note that relaxing the constraint to allow $\mathrm{comp}\circ\mathrm{rev}$ pairings yields a lower-cost matching (96 vs 120), but requires per-orbit case analysis rather than the uniform reverse-priority rule.
1.1 Main Results
Theorem 1.1 (Uniqueness of Optimal $\mathrm{comp}/\mathrm{rev}$ Matching).
For $n=6$, there exists a unique $K_{4}$-equivariant perfect matching on $\{0,1\}^{6}$ minimizing total Hamming distance among matchings where each pair is either $\{h,\mathrm{comp}(h)\}$ or $\{h,\mathrm{rev}(h)\}$. This matching is given by the reverse-priority rule:
$\mathrm{partner}(h)=\begin{cases}\mathrm{rev}(h)&\text{if }h\neq\mathrm{rev}(h)\\ \mathrm{comp}(h)&\text{if }h=\mathrm{rev}(h)\end{cases}$
The total Hamming cost is $120$, compared to $192$ for the complement-only matching.
Theorem 1.2 (Isomorphism with King Wen).
The King Wen sequence of the I Ching defines a perfect matching on $\{0,1\}^{6}$ that is isomorphic to the unique optimal $K_{4}$-equivariant matching of Theorem 1.1.
1.2 Formal Verification
All theorems are formally verified in Lean 4 with the Mathlib library. Key modules:
  • IChing/Hexagram.lean: $\{0,1\}^{6}$ representation, Hamming distance
  • IChing/Symmetry.lean: $K_{4}$-action, orbit structure
  • IChing/KingWenOptimality.lean: Optimality and uniqueness proofs
2 Preliminaries
2.1 Hexagrams as Binary Vectors
Definition 2.1 (Hexagram).
A hexagram is an element $h\in\{0,1\}^{6}=\{0,1\}^{6}$. We index positions 0 (bottom) through $5$ (top), following the traditional convention that hexagrams are read from bottom to top. The correspondence between hexagrams and binary numbers was noted by Leibniz [2].
There are $2^{6}=64$ hexagrams, corresponding to the vertices of the 6-dimensional hypercube.
Definition 2.2 (Hamming Distance).
For $h_{1},h_{2}\in\{0,1\}^{6}$, the Hamming distance is
$d_{H}(h_{1},h_{2})=\#\{i:h_{1}(i)\neq h_{2}(i)\}=\sum_{i=0}^{5}|h_{1}(i)-h_{2}(i)|$
Definition 2.3 (Total Hamming Cost of a Matching).
For a perfect matching $M$ on $\{0,1\}^{6}$ (a partition into 32 disjoint pairs), the total Hamming cost is the sum of Hamming distances over all pairs:
$\mathrm{Cost}(M)=\sum_{\{h_{1},h_{2}\}\in M}d_{H}(h_{1},h_{2})$
This measures how “different” the paired elements are overall.
2.2 The Klein Four-Group Action
Definition 2.4 (Complement and Reversal).
Define operations on hexagrams:
$\mathrm{comp}(h)(i)$ $=1-h(i)\quad\text{(bitwise complement)}$
$\mathrm{rev}(h)(i)$ $=h(5-i)\quad\text{(bit reversal)}$
Proposition 2.5.
The operations $\mathrm{comp}$ and $\mathrm{rev}$ satisfy:
  1. 1.
    $\mathrm{comp}\circ\mathrm{comp}=\mathrm{id}$ (complement is an involution)
  2. 2.
    $\mathrm{rev}\circ\mathrm{rev}=\mathrm{id}$ (reversal is an involution)
  3. 3.
    $\mathrm{comp}\circ\mathrm{rev}=\mathrm{rev}\circ\mathrm{comp}$ (they commute)
Thus $\{\mathrm{id},\mathrm{comp},\mathrm{rev},\mathrm{comp}\circ\mathrm{rev}\}$ forms the Klein four-group $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$.
Definition 2.6 (Orbit).
The orbit of a hexagram $h$ under the Klein four-group is
$\mathrm{orbit}(h)=\{h,\mathrm{comp}(h),\mathrm{rev}(h),\mathrm{comp}(\mathrm{rev}(h))\}$
Proposition 2.7 (Orbit Sizes).
For any hexagram $h$, $|\mathrm{orbit}(h)|\in\{2,4\}$.
  • Size 4: generic hexagrams (48 total, forming 12 orbits)
  • Size 2: palindromes with $h=\mathrm{rev}(h)$ (8 total, forming 4 orbits)
  • Size 2: anti-symmetric hexagrams with $\mathrm{rev}(h)=\mathrm{comp}(h)$ but $h\neq\mathrm{rev}(h)$ (8 total, forming 4 orbits)
Definition 2.8 (Anti-Symmetric Hexagram).
A hexagram $h$ is anti-symmetric if $h(i)\neq h(5-i)$ for all $i\in\{0,1,2\}$. Equivalently, $\mathrm{rev}(h)=\mathrm{comp}(h)$.
Definition 2.9 (Palindrome).
A hexagram $h$ is a palindrome if $\mathrm{rev}(h)=h$.
2.3 Hamming Distances of Group Actions
Proposition 2.10.
For any hexagram $h$:
  1. 1.
    $d_{H}(h,\mathrm{comp}(h))=6$ (complement flips all bits)
  2. 2.
    $d_{H}(h,\mathrm{rev}(h))\leq 6$, with equality iff $h(i)\neq h(5-i)$ for all $i$
Lemma 2.11 (Key Lemma).
If $d_{H}(h,\mathrm{rev}(h))=6$, then $\mathrm{rev}(h)=\mathrm{comp}(h)$.
Proof.
If all 6 positions differ between $h$ and $\mathrm{rev}(h)$, then for each $i$: $\mathrm{rev}(h)(i)=h(5-i)\neq h(i)$. Since bits are in $\{0,1\}$, this means $h(5-i)=1-h(i)=\mathrm{comp}(h)(i)$. Thus $\mathrm{rev}(h)=\mathrm{comp}(h)$. ∎
This lemma is crucial: when reversal achieves maximum distance 6, it coincides with complement, so there is no choice to make.
3 The King Wen Sequence
3.1 Definition
The King Wen sequence assigns each hexagram a number 1–64 according to tradition [1]. We represent it via a mapping from King Wen numbers to binary:
Definition 3.1 (King Wen Binary Mapping).
Define $\mathrm{KW}:\{0,\ldots,63\}\to\{0,1\}^{6}$ as follows (index 0 corresponds to King Wen number 1):
KW 1 2 3 4 5 6 7 8
Binary 63 0 17 34 23 58 2 16
KW 9 10 11 12 13 14 15 16
Binary 55 59 7 56 61 47 4 8
(Full table in Appendix.)
Definition 3.2 (King Wen Pairs).
The King Wen pairs are $(h_{2k},h_{2k+1})$ for $k=0,\ldots,31$, where $h_{i}=\mathrm{KW}(i)$.
3.2 The Equivariance Theorem
Theorem 3.3 (100% Structural Regularity).
Every King Wen pair $(h_{1},h_{2})$ satisfies:
$h_{2}=\mathrm{comp}(h_{1})\quad\text{or}\quad h_{2}=\mathrm{rev}(h_{1})$
The 32 pairs partition as:
  • 4 pairs: palindromes, paired by complement ($h_{2}=\mathrm{comp}(h_{1})$, distance 6)
  • 4 pairs: anti-symmetric, paired by reversal == complement ($h_{2}=\mathrm{rev}(h_{1})=\mathrm{comp}(h_{1})$, distance 6)
  • 24 pairs: generic, paired by reversal ($h_{2}=\mathrm{rev}(h_{1})$, distance 2 or 4)
Proof.
Verified computationally over all 32 pairs. In Lean 4: decide. ∎
Corollary 3.4.
The King Wen pairing respects the Klein four-group: paired hexagrams always lie in the same orbit.
4 Optimality of the King Wen Matching
4.1 The Reverse-Priority Rule
Definition 4.1 (Reverse-Priority Matching).
Define the priority partner function:
$\mathrm{partner}(h)=\begin{cases}\mathrm{comp}(h)&\text{if }h=\mathrm{rev}(h)\text{ (palindrome)}\\ \mathrm{rev}(h)&\text{otherwise}\end{cases}$
The intuition: prefer reversal (which has distance $\leq 6$) unless reversal is trivial (palindrome), in which case use complement.
Theorem 4.2 (Involution).
The priority partner function is an involution: $\mathrm{partner}(\mathrm{partner}(h))=h$.
Proof.
Two cases:
  1. 1.
    If $h$ is a palindrome, then $\mathrm{comp}(h)$ is also a palindrome (complement preserves palindrome structure), so $\mathrm{partner}(\mathrm{comp}(h))=\mathrm{comp}(\mathrm{comp}(h))=h$.
  2. 2.
    If $h$ is not a palindrome, then $\mathrm{rev}(h)$ is also not a palindrome, so $\mathrm{partner}(\mathrm{rev}(h))=\mathrm{rev}(\mathrm{rev}(h))=h$.
4.2 Cost Minimization
Definition 4.3 (Total Hamming Cost).
For a perfect matching $M$ on $\{0,1\}^{6}$, the total Hamming cost is
$\mathrm{Cost}(M)=\sum_{\{h_{1},h_{2}\}\in M}d_{H}(h_{1},h_{2})$
Theorem 4.4 (Optimality among $\mathrm{comp}/\mathrm{rev}$ Matchings).
Among $K_{4}$-equivariant perfect matchings on $\{0,1\}^{6}$ where each pair uses $\mathrm{comp}$ or $\mathrm{rev}$, the reverse-priority matching uniquely minimizes total Hamming cost.
Proof.
Restricting to $\mathrm{comp}/\mathrm{rev}$ pairings, the options for each hexagram $h$ are $\mathrm{comp}(h)$ or $\mathrm{rev}(h)$.
  • By Proposition 2.10, $d_{H}(h,\mathrm{rev}(h))\leq 6=d_{H}(h,\mathrm{comp}(h))$.
  • Equality holds iff $\mathrm{rev}(h)=\mathrm{comp}(h)$ (Lemma 2.11).
  • For palindromes, $\mathrm{rev}(h)=h$, so the only non-trivial option is $\mathrm{comp}(h)$.
Thus the reverse-priority rule chooses the minimum-distance option at each step. Since there is no actual choice when distances are equal (the options coincide), the matching is uniquely determined. ∎
Remark 4.5 (The $\mathrm{comp}\circ\mathrm{rev}$ Alternative).
On a size-4 orbit, a third equivariant pairing exists: $\{h,\mathrm{comp}(\mathrm{rev}(h))\}$. For orbits where $d_{H}(h,\mathrm{rev}(h))=4$, we have $d_{H}(h,\mathrm{comp}(\mathrm{rev}(h)))=2$, making the $\mathrm{comp}\circ\mathrm{rev}$ pairing strictly cheaper. A mixed strategy using $\mathrm{comp}\circ\mathrm{rev}$ on such orbits achieves total cost $96<120$. However, this requires case-by-case analysis of each orbit, whereas the reverse-priority rule is a uniform principle. The King Wen sequence follows the uniform rule rather than the per-orbit optimum.
Corollary 4.6 (Cost Values).
The total Hamming cost of the reverse-priority matching:
  • Palindrome pairs: $4\times 6=24$
  • Anti-symmetric pairs: $4\times 6=24$
  • Generic pairs: $12\times 2+12\times 4=72$
Total: $24+24+72=120$.
For comparison, the complement-only matching has cost $32\times 6=192$.
4.3 Uniqueness
Theorem 4.7 (Uniqueness).
Any equivariant matching satisfying the reverse-priority rule equals the priority partner function.
Proof.
The reverse-priority rule uniquely specifies the partner at each hexagram. By Theorem 4.2, this defines a valid involution, hence a perfect matching. ∎
Corollary 4.8 (King Wen is Canonical).
The King Wen sequence encodes the unique cost-minimizing $\mathrm{comp}/\mathrm{rev}$ equivariant matching on $\{0,1\}^{6}$.
5 Discussion
5.1 Orbit Structure
The $K_{4}$-action on $\{0,1\}^{6}$ partitions the 64 elements into orbits of size 2 or 4:
  • Size-4 orbits: 12 orbits containing 48 generic elements
  • Size-2 orbits: 8 orbits containing 16 palindromes (elements fixed by $\mathrm{rev}$)
Within each orbit, the equivariant matching must pair elements. For size-4 orbits, there are three equivariant pairings induced by the non-identity involutions $\mathrm{comp}$, $\mathrm{rev}$, and $\mathrm{comp}\circ\mathrm{rev}$. The reverse-priority rule uses $\mathrm{rev}$ (falling back to $\mathrm{comp}$ for palindromes), which minimizes cost among $\mathrm{comp}/\mathrm{rev}$ matchings. A mixed strategy using $\mathrm{comp}\circ\mathrm{rev}$ on orbits where $d_{H}(h,\mathrm{rev}(h))=4$ achieves the global minimum cost of 96.
5.2 Uniqueness Mechanism
The key to uniqueness is Lemma 2.11: when $d_{H}(h,\mathrm{rev}(h))=6$, we have $\mathrm{rev}(h)=\mathrm{comp}(h)$, so the two candidate partners coincide. This eliminates all apparent degrees of freedom in the optimization.
5.3 Formal Verification
All results are machine-checked in Lean 4 using the Mathlib library:
  • decide tactic for computational verification of all 32 pairs
  • Constructive proof that the reverse-priority function is an involution
  • Proof of optimality via case analysis on orbit structure
5.4 Subgroup Matchings
One may ask whether the theorem holds for subgroups of $K_{4}$:
  • Complement-only ($\langle\mathrm{comp}\rangle$): Valid perfect matching with cost 192, trivially unique since each hexagram has exactly one complement.
  • Reversal-only ($\langle\mathrm{rev}\rangle$): Not a valid perfect matching. The 8 palindromes satisfy $\mathrm{rev}(h)=h$, so they cannot pair with their own reversal. A perfect matching requires pairing with a different element.
Thus the reverse-priority rule is the minimal hybrid that produces a valid matching: use $\mathrm{rev}$ whenever possible, fall back to $\mathrm{comp}$ for the 8 palindromes.
5.5 Extensions
The $K_{4}$-action generalizes to $\{0,1\}^{n}$ for any $n$. The orbit structure and optimal matching problem remain well-defined; we conjecture the reverse-priority rule remains optimal for all $n$.
Appendix A Full King Wen Binary Table
KW Bin KW Bin KW Bin KW Bin
1 63 17 25 33 60 49 29
2 0 18 38 34 15 50 46
3 17 19 3 35 40 51 9
4 34 20 48 36 5 52 36
5 23 21 41 37 53 53 52
6 58 22 37 38 43 54 11
7 2 23 32 39 20 55 13
8 16 24 1 40 10 56 44
9 55 25 57 41 35 57 54
10 59 26 39 42 49 58 27
11 7 27 33 43 31 59 50
12 56 28 30 44 62 60 19
13 61 29 18 45 24 61 51
14 47 30 45 46 6 62 12
15 4 31 28 47 26 63 21
16 8 32 14 48 22 64 42
Binary values represent the hexagram as a 6-bit integer, with bit 0 as the bottom line.