### v2

# Optimal Equivariant Matchings on the 6-Cube With an Application to the King Wen Sequence

Abstract
We characterize perfect matchings on the Boolean hypercube $\{0,1\}^{n}$ that are equivariant under the Klein four-group $K_{4}=\langle\mathrm{comp},\mathrm{rev}\rangle$ generated by bitwise complement and reversal. For $n=6$, we prove there exists a unique $K_{4}$-equivariant matching minimizing total Hamming cost among matchings using only $\mathrm{comp}$ or $\mathrm{rev}$ pairings, achieving cost 120 versus 192 for the complement-only matching. The optimal matching is determined by a simple “reverse-priority rule”: pair each element with its reversal unless it is a palindrome, in which case pair with its complement. We verify that the historically significant King Wen sequence of the I Ching is isomorphic to this optimal matching. Notably, allowing $\mathrm{comp}\circ\mathrm{rev}$ pairings yields lower cost (96), but the King Wen sequence follows the structurally simpler rule. All results are formally verified in Lean 4 with the Mathlib library.
1 Introduction
The Boolean hypercube $\{0,1\}^{n}=\{0,1\}^{n}$ admits two natural involutions: bitwise complement $\mathrm{comp}$, which flips each bit ($0\leftrightarrow 1$), and bit reversal $\mathrm{rev}$, which reverses bit order ($b_{0}b_{1}\cdots b_{n-1}\mapsto b_{n-1}\cdots b_{1}b_{0}$). Together these generate the Klein four-group $K_{4}=\{\mathrm{id},\mathrm{comp},\mathrm{rev},\mathrm{comp}\circ\mathrm{rev}\}$.
A $\mathrm{comp}/\mathrm{rev}$ matching is a perfect matching where each pair $\{h,h^{\prime}\}$ satisfies either $h^{\prime}=\mathrm{comp}(h)$ or $h^{\prime}=\mathrm{rev}(h)$. More generally, a $K_{4}$-equivariant matching also permits $h^{\prime}=\mathrm{comp}(\mathrm{rev}(h))$.
This paper addresses the optimization problem:
Among $\mathrm{comp}/\mathrm{rev}$ matchings on $\{0,1\}^{n}$, which minimize total Hamming distance (Definition 2.4)?
For $n=6$, one might expect this to be trivial: just pair each element with the cheaper option. But in most equivariant matching problems, locally optimal choices conflict globally—choosing the cheapest partner for $h$ may force a suboptimal choice elsewhere. Our main result is that for $\{0,1\}^{6}$, no such conflicts arise: a simple greedy rule yields the unique global optimum. The key is Lemma 2.12, which shows that $\mathrm{comp}$ and $\mathrm{rev}$ never genuinely compete—whenever they differ in cost, $\mathrm{rev}$ is strictly cheaper.
As an application, we verify that the King Wen sequence—a traditional ordering of the 64 hexagrams of the I Ching into 32 pairs—is isomorphic to this unique optimal matching. We also note that relaxing to full $K_{4}$-equivariant matchings (allowing $\mathrm{comp}\circ\mathrm{rev}$ pairings) yields lower cost (96 vs 120), but requires per-orbit case analysis rather than a uniform rule.
1.1 Main Results
Theorem 1.1 (Uniqueness of Optimal $\mathrm{comp}/\mathrm{rev}$ Matching).
For $n=6$, there exists a unique $K_{4}$-equivariant perfect matching on $\{0,1\}^{6}$ minimizing total Hamming distance among matchings where each pair is either $\{h,\mathrm{comp}(h)\}$ or $\{h,\mathrm{rev}(h)\}$. This matching is given by the reverse-priority rule:
$\mathrm{partner}(h)=\begin{cases}\mathrm{rev}(h)&\text{if }h\neq\mathrm{rev}(h)\\ \mathrm{comp}(h)&\text{if }h=\mathrm{rev}(h)\end{cases}$
The total Hamming cost is $120$, compared to $192$ for the complement-only matching.
Theorem 1.2 (Isomorphism with King Wen).
The King Wen sequence of the I Ching defines a perfect matching on $\{0,1\}^{6}$ that is isomorphic to the unique optimal $K_{4}$-equivariant matching of Theorem $1.$
1.2 Formal Verification
All theorems are formally verified in Lean 4 with the Mathlib library. Key modules:
  • IChing/Hexagram.lean: $\{0,1\}^{6}$ representation, Hamming distance
  • IChing/Symmetry.lean: $K_{4}$-action, orbit structure
  • IChing/KingWenOptimality.lean: Optimality and uniqueness proofs
2 Preliminaries
2.1 Hexagrams as Binary Vectors
In the I Ching, a hexagram is a figure composed of six stacked horizontal lines, each either solid (representing yang) or broken (representing yin). Reading solid as 11 and broken as 0, from bottom to top, each hexagram corresponds to a 6-bit binary string. There are $2^{6}=64$ hexagrams, corresponding to the vertices of the 6-dimensional hypercube.
Definition 2.1 (Hexagram).
A hexagram is an element $h\in\{0,1\}^{6}$, with positions indexed 0 (bottom) through 55 (top).
Remark $2.$
The correspondence between hexagrams and binary numbers was noted by Leibniz [2].
Definition 2.3 (Hamming Distance).
For $h_{1},h_{2}\in\{0,1\}^{6}$, the Hamming distance is
$d_{H}(h_{1},h_{2})=\#\{i:h_{1}(i)\neq h_{2}(i)\}=\sum_{i=0}^{5}|h_{1}(i)-h_{2}(i)|$
Definition 2.4 (Total Hamming Cost of a Matching).
For a perfect matching $M$ on $\{0,1\}^{6}$ (a partition into 32 disjoint pairs), the total Hamming cost is the sum of Hamming distances over all pairs:
$\mathrm{Cost}(M)=\sum_{\{h_{1},h_{2}\}\in M}d_{H}(h_{1},h_{2})$
This measures how “different” the paired elements are overall.
2.2 The Klein Four-Group Action
Definition 2.5 (Complement and Reversal).
Define operations on hexagrams:
$\mathrm{comp}(h)(i)$ $=1-h(i)\quad\text{(bitwise complement)}$
$\mathrm{rev}(h)(i)$ $=h(5-i)\quad\text{(bit reversal)}$
Proposition 2.6.
The operations $\mathrm{comp}$ and $\mathrm{rev}$ satisfy:
  1. 1.
    $\mathrm{comp}\circ\mathrm{comp}=\mathrm{id}$ (complement is an involution)
  2. 2.
    $\mathrm{rev}\circ\mathrm{rev}=\mathrm{id}$ (reversal is an involution)
  3. 3.
    $\mathrm{comp}\circ\mathrm{rev}=\mathrm{rev}\circ\mathrm{comp}$ (they commute)
Thus $\{\mathrm{id},\mathrm{comp},\mathrm{rev},\mathrm{comp}\circ\mathrm{rev}\}$ forms the Klein four-group $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$.
Definition 2.7 (Orbit).
The orbit of a hexagram $h$ under the Klein four-group is
$\mathrm{orbit}(h)=\{h,\mathrm{comp}(h),\mathrm{rev}(h),\mathrm{comp}(\mathrm{rev}(h))\}$
Proposition 2.8 (Orbit Sizes).
For any hexagram $h$, $|\mathrm{orbit}(h)|\in\{2,4\}$ (see Figure 1).
  • Size 4: generic hexagrams (48 total, forming 12 orbits)
  • Size 2: palindromes with $h=\mathrm{rev}(h)$ (8 total, forming 4 orbits)
  • Size 2: anti-symmetric hexagrams with $\mathrm{rev}(h)=\mathrm{comp}(h)$ but $h\neq\mathrm{rev}(h)$ (8 total, forming 4 orbits)
[…]
Figure 1: Orbit types under $K_{4}$. Edge labels show Hamming distances. In (a), the $\mathrm{rev}$ edge (blue) has cost 2244, always $\leq$ the $\mathrm{comp}$ edge (cost 66). In (b) and (c), only one non-trivial pairing exists.
Definition 2.9 (Anti-Symmetric Hexagram).
A hexagram $h$ is anti-symmetric if $h(i)\neq h(5-i)$ for all $i\in\{0,1,2\}$. Equivalently, $\mathrm{rev}(h)=\mathrm{comp}(h)$.
Definition 2.10 (Palindrome).
A hexagram $h$ is a palindrome if $\mathrm{rev}(h)=h$.
2.3 Hamming Distances of Group Actions
Proposition 2.11.
For any hexagram $h$:
  1. 1.
    $d_{H}(h,\mathrm{comp}(h))=6$ (complement flips all bits)
  2. 2.
    $d_{H}(h,\mathrm{rev}(h))\leq 6$, with equality iff $h(i)\neq h(5-i)$ for all $i$
Lemma 2.12 (No-Conflict Lemma).
For any hexagram $h$, exactly one of the following holds:
  1. 1.
    $d_{H}(h,\mathrm{rev}(h))<d_{H}(h,\mathrm{comp}(h))=6$ (reversal is strictly cheaper), or
  2. 2.
    $\mathrm{rev}(h)=\mathrm{comp}(h)$ and both distances equal 66 (the options coincide), or
  3. 3.
    $\mathrm{rev}(h)=h$ (palindrome, so $\mathrm{rev}$ yields trivial pairing).
In particular, $\mathrm{rev}$ and $\mathrm{comp}$ never genuinely compete: when both are valid and distinct, $\mathrm{rev}$ is strictly cheaper.
Proof.
Since $\mathrm{comp}$ flips all bits, $d_{H}(h,\mathrm{comp}(h))=6$ always. For $\mathrm{rev}$: if $d_{H}(h,\mathrm{rev}(h))=6$, then every position differs, so $h(5-i)\neq h(i)$ for all ii. This means $h(5-i)=1-h(i)=\mathrm{comp}(h)(i)$, hence $\mathrm{rev}(h)=\mathrm{comp}(h)$. If $\mathrm{rev}(h)=h$, then $h$ is a palindrome. Otherwise $d_{H}(h,\mathrm{rev}(h))<6$. ∎
Remark 2.13 (Why Greedy Works).
In typical equivariant matching problems, choosing the locally optimal partner for $h$ may force suboptimal choices elsewhere. The No-Conflict Lemma eliminates this tension: the greedy choice (prefer $\mathrm{rev}$ unless it’s trivial) is unambiguous. There is no trade-off to optimize, hence no room for conflict (see Figure 2). This is why the simple reverse-priority rule achieves the unique global optimum.
[…]
Figure 2: Forced consistency of the reverse-priority rule. Choosing $\mathrm{rev}(h)$ as partner for $h$ simultaneously determines $h$ as partner for $\mathrm{rev}(h)$, since $\mathrm{rev}$ is an involution. No secondary decision can conflict.
3 The King Wen Sequence
3.1 Definition
The King Wen sequence assigns each hexagram a number 1–64 according to tradition [1]. We represent it via a mapping from King Wen numbers to binary:
Definition 3.1 (King Wen Binary Mapping).
Define $\mathrm{KW}:\{0,\ldots,63\}\to\{0,1\}^{6}$ as follows (index 0 corresponds to King Wen number 1):
KW 1 2 3 4 5 6 7 8
Binary 63 0 17 34 23 58 2 16
KW 9 10 11 12 13 14 15 16
Binary 55 59 7 56 61 47 4 8
(Full table in Appendix.)
Definition 3.2 (King Wen Pairs).
The King Wen pairs are $(h_{2k},h_{2k+1})$ for $k=0,\ldots,31$, where $h_{i}=\mathrm{KW}(i)$.
3.2 The Equivariance Theorem
Theorem 3.3 (100% Structural Regularity).
Every King Wen pair $(h_{1},h_{2})$ satisfies:
$h_{2}=\mathrm{comp}(h_{1})\quad\text{or}\quad h_{2}=\mathrm{rev}(h_{1})$
The 32 pairs partition as:
  • 4 pairs: palindromes, paired by complement ($h_{2}=\mathrm{comp}(h_{1})$, distance 6)
  • 4 pairs: anti-symmetric, paired by reversal == complement ($h_{2}=\mathrm{rev}(h_{1})=\mathrm{comp}(h_{1})$, distance 6)
  • 24 pairs: generic, paired by reversal ($h_{2}=\mathrm{rev}(h_{1})$, distance 2 or 4)
Proof.
Verified computationally over all 32 pairs. In Lean 4: decide. ∎
Corollary 3.4.
The King Wen pairing respects the Klein four-group: paired hexagrams always lie in the same orbit.
4 Optimality of the King Wen Matching
4.1 The Reverse-Priority Rule
Definition 4.1 (Reverse-Priority Matching).
Define the priority partner function:
$\mathrm{partner}(h)=\begin{cases}\mathrm{comp}(h)&\text{if }h=\mathrm{rev}(h)\text{ (palindrome)}\\ \mathrm{rev}(h)&\text{otherwise}\end{cases}$
The intuition: prefer reversal (which has distance $\leq 6$) unless reversal is trivial (palindrome), in which case use complement.
Theorem 4.2 (Involution).
The priority partner function is an involution: $\mathrm{partner}(\mathrm{partner}(h))=h$.
Proof.
Two cases:
  1. 1.
    If $h$ is a palindrome, then $\mathrm{comp}(h)$ is also a palindrome (complement preserves palindrome structure), so $\mathrm{partner}(\mathrm{comp}(h))=\mathrm{comp}(\mathrm{comp}(h))=h$.
  2. 2.
    If $h$ is not a palindrome, then $\mathrm{rev}(h)$ is also not a palindrome, so $\mathrm{partner}(\mathrm{rev}(h))=\mathrm{rev}(\mathrm{rev}(h))=h$.
4.2 Cost Minimization
Definition 4.3 (Total Hamming Cost).
For a perfect matching $M$ on $\{0,1\}^{6}$, the total Hamming cost is
$\mathrm{Cost}(M)=\sum_{\{h_{1},h_{2}\}\in M}d_{H}(h_{1},h_{2})$
Theorem 4.4 (Optimality among $\mathrm{comp}/\mathrm{rev}$ Matchings).
Among $K_{4}$-equivariant perfect matchings on $\{0,1\}^{6}$ where each pair uses $\mathrm{comp}$ or $\mathrm{rev}$, the reverse-priority matching uniquely minimizes total Hamming cost.
Proof.
Restricting to $\mathrm{comp}/\mathrm{rev}$ pairings, the options for each hexagram $h$ are $\mathrm{comp}(h)$ or $\mathrm{rev}(h)$.
  • By Proposition 2.11, $d_{H}(h,\mathrm{rev}(h))\leq 6=d_{H}(h,\mathrm{comp}(h))$.
  • Equality holds iff $\mathrm{rev}(h)=\mathrm{comp}(h)$ (Lemma 2.12).
  • For palindromes, $\mathrm{rev}(h)=h$, so the only non-trivial option is $\mathrm{comp}(h)$.
Thus the reverse-priority rule chooses the minimum-distance option at each step. Since there is no actual choice when distances are equal (the options coincide), the matching is uniquely determined. ∎
Remark 4.5 (The $\mathrm{comp}\circ\mathrm{rev}$ Alternative).
On a size-4 orbit, a third equivariant pairing exists: $\{h,\mathrm{comp}(\mathrm{rev}(h))\}$. For orbits where $d_{H}(h,\mathrm{rev}(h))=4$, we have $d_{H}(h,\mathrm{comp}(\mathrm{rev}(h)))=2$, making the $\mathrm{comp}\circ\mathrm{rev}$ pairing strictly cheaper (see Figure 3). A mixed strategy using $\mathrm{comp}\circ\mathrm{rev}$ on such orbits achieves total cost $96<120$. However, this requires case-by-case analysis of each orbit, whereas the reverse-priority rule is a uniform principle. The King Wen sequence follows the uniform rule rather than the per-orbit optimum.
[…]
Figure 3: A size-4 orbit where $\mathrm{comp}\circ\mathrm{rev}$ is optimal. The diagonal pairing (dashed) has cost 22, versus 44 for reversal (solid). Using $\mathrm{comp}\circ\mathrm{rev}$ on such orbits reduces total cost from $120$ to $96$, but requires per-orbit analysis.
Corollary 4.6 (Cost Values).
The total Hamming cost of the reverse-priority matching:
  • Palindrome pairs: $4\times 6=24$
  • Anti-symmetric pairs: $4\times 6=24$
  • Generic pairs: $12\times 2+12\times 4=72$
Total: $24+24+72=120$.
For comparison, the complement-only matching has cost $32\times 6=192$.
4.3 Uniqueness
Theorem 4.7 (Uniqueness).
Any equivariant matching satisfying the reverse-priority rule equals the priority partner function.
Proof.
The reverse-priority rule uniquely specifies the partner at each hexagram. By Theorem 4.2, this defines a valid involution, hence a perfect matching. ∎
Corollary 4.8 (King Wen is Canonical).
The King Wen sequence encodes the unique cost-minimizing $\mathrm{comp}/\mathrm{rev}$ equivariant matching on $\{0,1\}^{6}$.
5 Discussion
5.1 Orbit Structure
The $K_{4}$-action on $\{0,1\}^{6}$ partitions the 64 elements into orbits of size 2 or 4:
  • Size-4 orbits: 12 orbits containing 48 generic elements
  • Size-2 orbits: 8 orbits containing 16 palindromes (elements fixed by $\mathrm{rev}$)
Within each orbit, the equivariant matching must pair elements. For size-4 orbits, there are three equivariant pairings induced by the non-identity involutions $\mathrm{comp}$, $\mathrm{rev}$, and $\mathrm{comp}\circ\mathrm{rev}$. The reverse-priority rule uses $\mathrm{rev}$ (falling back to $\mathrm{comp}$ for palindromes), which minimizes cost among $\mathrm{comp}/\mathrm{rev}$ matchings. A mixed strategy using $\mathrm{comp}\circ\mathrm{rev}$ on orbits where $d_{H}(h,\mathrm{rev}(h))=4$ achieves the global minimum cost of 96.
5.2 Uniqueness Mechanism
The key to uniqueness is the No-Conflict Lemma 2.12: $\mathrm{comp}$ and $\mathrm{rev}$ never genuinely compete. When they have different costs, $\mathrm{rev}$ wins; when they have equal cost, they coincide. This eliminates all apparent degrees of freedom, making greedy optimization globally consistent.
5.3 Formal Verification
All results are machine-checked in Lean 4 using the Mathlib library:
  • decide tactic for computational verification of all 32 pairs
  • Constructive proof that the reverse-priority function is an involution
  • Proof of optimality via case analysis on orbit structure
5.4 Subgroup Matchings
One may ask whether the theorem holds for subgroups of $K_{4}$:
  • Complement-only ($\langle\mathrm{comp}\rangle$): Valid perfect matching with cost 192, trivially unique since each hexagram has exactly one complement.
  • Reversal-only ($\langle\mathrm{rev}\rangle$): Not a valid perfect matching. The 8 palindromes satisfy $\mathrm{rev}(h)=h$, so they cannot pair with their own reversal. A perfect matching requires pairing with a different element.
Thus the reverse-priority rule is the minimal hybrid that produces a valid matching: use $\mathrm{rev}$ whenever possible, fall back to $\mathrm{comp}$ for the 8 palindromes.
5.5 Extensions
The $K_{4}$-action generalizes to $\{0,1\}^{n}$ for any $n$. The orbit structure and optimal matching problem remain well-defined; we conjecture the reverse-priority rule remains optimal for all $n$.
Appendix A Full King Wen Binary Table
KW Bin KW Bin KW Bin KW Bin
1 63 17 25 33 60 49 29
2 0 18 38 34 15 50 46
3 17 19 3 35 40 51 9
4 34 20 48 36 5 52 36
5 23 21 41 37 53 53 52
6 58 22 37 38 43 54 11
7 2 23 32 39 20 55 13
8 16 24 1 40 10 56 44
9 55 25 57 41 35 57 54
10 59 26 39 42 49 58 27
11 7 27 33 43 31 59 50
12 56 28 30 44 62 60 19
13 61 29 18 45 24 61 51
14 47 30 45 46 6 62 12
15 4 31 28 47 26 63 21
16 8 32 14 48 22 64 42
Binary values represent the hexagram as a 6-bit integer, with bit 0 as the bottom line.